OpenStudy (anonymous):

FAN AND MEDAL PLS Triangle ABC is congruent to triangle DEF. In triangle ABC, side AB measures 13, side BC measures x+13, and side CA measures 16. In triangle DEF, side DE measures 13, side EF measures 3x−20, and side FD measures 16. What equation would help you to solve for the side length of BC and EF? Explain your reasoning using complete sentences.

OpenStudy (anonymous):

@jim_thompson5910 @dan815 @confluxepic @Loser66 @TheSmartOne @Whitemonsterbunny17 @eliassaab @bibby @EclipsedStar @sleepyjess @bohotness

OpenStudy (anonymous):

x+13 = 3x-20 is the equation, I just need help on explaining the reasoning using complete sentences.

OpenStudy (anonymous):

well heres what your triangles will look like, just to give you an idea 

OpenStudy (anonymous):

I got that, thank you. can you just help me with the rest

OpenStudy (anonymous):

@kittenlover731

OpenStudy (anonymous):

Reorder the terms: Note: you can or can not do this if you want 13 + x = -20 + 3x

OpenStudy (anonymous):

subtract 13 from both sides 13 + x -13= -20 + 3x -13

OpenStudy (anonymous):

That doesnt answer the second question, that just simplify it.

OpenStudy (anonymous):

@YinYangInsane